40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 1 14:56:08 2012, Jamie, Update, SUS, Oplevs all different, messed up oplevs.png
    Reply  Fri Jun 1 18:07:32 2012, steve, Update, SUS, Oplevs servo values 
    Reply  Fri Jun 1 19:48:54 2012, rana, Update, SUS, Oplevs all different, messed up 
    Reply  Wed Jul 11 16:19:08 2012, yuta, Update, SUS, oplev servo phiology 
Message ID: 6964     Entry time: Wed Jul 11 16:19:08 2012     In reply to: 6743
Author: yuta 
Type: Update 
Category: SUS 
Subject: oplev servo phiology 

I heard that Steve did great work on oplev in Feb 2012.
Here's summary what happened to oplev since then.
Someone changed oplev filters and gain. I couldn't find elog about it. Does anyone know?

Quadrant sum:
  Quadrant sum (C1:SUS-XXX_OLSUM) for each optic now and in Feb 2012 is

  ITMX   ITMY   ETMX   ETMY     BS    PRM    SRM
  2456  14630   1476  14885   3650   4302   2937   counts (now)
  1300  14500    900   9000   3500   4000   2600   counts (Feb 6, 2012 elog
#6256)
 0.025  0.3    0.2    0.2    0.05   0.06   0.04    mW on QPD (Feb 6, 2012 elog
#6256)
  1350  15000   1500  15500   3500   4000   2600   counts (Feb 23, 2012 elog
#6744)

  ETMX oplev laser was replaced on May 22, 2012, and quadrant sum was 20500 counts at that time (elog #6656).


Oplev servo openloop transfer functions:
  In Feb 2012, gains were adjusted and filter settings are recorded by Steve.
  For all pitch OLTF, see elog #6309.
  For all yaw OLTF, see elog #6323.

  All the filters in Feb is listed in elog #6744.
  Filters now are messed up, as Jamie pointed out in elog #6743.
  Below is the current filter settings.

  I turned ELP and RLP filters on, which wasn't on to cut-off noises at higher frequencies.
  I left resonant gains of ETMs because I don't know what they are for.
  I put ELP35 for ITMs, BS, PRM and SRM. I put RLP80 for BS, PRM and SRM.
  I will leave ELP35 off for BS and SRM because they oscillate currently. ELP50 and ELP40 is on for a substitution. I will readjust them soon.

  I don't know who changed all gains (except for PRM, which I adjusted in elog #6952). It doen't look like it is because of change in quadrant sum.
  I also don't know who deleted 3.3 Hz resonant gain for BS.

  I put all similar filters in same place to make it organized. Now, basic fitlers are organized. We may need some resonant gains for each optics.

OPLEV SERVO 300^2:0 BR ELP RLP RES GAIN QPD counts
filter position FM1 FM5 FM9 FM4 FM3, FM4    
ETMY pit 300^2:0 BR 35 80

0.5 (off)

-0.2 (was -1.5) 14,900
ETMY yaw 300^2:0 BR 35 80 0.6 (off) -0.2 (was -1.0)  
ETMX pit 300^2:0 BR 35 80 0.5 (off) 0.5 1,500
ETMX yaw 300^2:0 BR 35 80 0.6 (off) 0.6 (was 1.0)  
ITMY pit 300^2:0 BR 35 80   2.1 (was 2.0) 14,600
ITMY yaw 300^2:0 BR 35 80   -2.0 (was -4.0)  
ITMX pit 300^2:0 BR 35 80   2.6 (was 1.0) 2,500
ITMX yaw 300^2:0 BR 35 80   -1.6 (was-2.0)  
BS pit 300^2:0 BR 50 (FM10) 80   0.6 (was 0.5) 3,700
BS yaw 300^2:0 BR 50 (FM10) 80 (3.3 is some how deleted) -0.6 (was -1.0)  
PRM pit 300^2:0 BR 35 80 3.3 (off) 0.15 (was 1.0) 4,300
PRM yaw 300^2:0 BR 35 80 3.3 (off),  4 (off) -0.2 (was 0.5)  
SRM pit 300^2:0 BR 40 (FM10) 80   -2.0 2,900
SRM yaw 300^2:0 BR 40 (FM10) 80   2.0  

  I also found Kiwamu's angular motion measurement during PRMI lock (elog #6320). They look different with my measurement yesterday (elog #6954).

ELOG V3.1.3-